# Researcher Population Pyramids for Tracking Global Demographic and Gender Trajectories

### v1

@@ Acknowledgments
Albania $8.98\%$ $-24.02\%$
Algeria $7.97\%$ $-39.88\%$
Australia $5.02\%$ $-30.00\%$
Austria $6.44\%$ $-40.30\%$
Bahrain $10.42\%$ $-39.85\%$
Belarus $9.03\%$ $-28.45\%$
Belgium $6.78\%$ $-36.45\%$
Bosnia and Herzegovina $6.23\%$ $-21.07\%$
Bulgaria $6.12\%$ $-4.91\%$
Canada $5.30\%$ $-32.75\%$
Cyprus $8.86\%$ $-45.05\%$
Czechia $5.28\%$ $-37.24\%$
Denmark $5.96\%$ $-40.05\%$
Egypt $9.83\%$ $-20.91\%$
Estonia $4.63\%$ $-39.35\%$
France $5.72\%$ $-30.46\%$
Germany $6.07\%$ $-44.18\%$
Ghana $11.84\%$ $-32.65\%$
Iceland $4.82\%$ $-30.83\%$
Iran, Islamic Republic of $7.43\%$ $-43.79\%$
Iraq $15.31\%$ $-24.69\%$
Ireland $5.59\%$ $-32.66\%$
Israel $6.18\%$ $-34.41\%$
Italy $6.85\%$ $-27.58\%$
Jamaica $4.71\%$ $-13.00\%$
Japan $4.98\%$ $-46.71\%$
Jordan $11.48\%$ $-34.67\%$
Kenya $9.08\%$ $-22.17\%$
Kuwait $7.81\%$ $-27.77\%$
Lebanon $8.71\%$ $-32.97\%$
Lithuania $5.86\%$ $-25.84\%$
Moldova, Republic of $13.65\%$ $-17.68\%$
Montenegro $6.05\%$ $-28.36\%$
Morocco $12.03\%$ $-33.47\%$
Netherlands $6.40\%$ $-36.12\%$
New Zealand $4.74\%$ $-34.23\%$
Nigeria $11.69\%$ $-16.68\%$
Norway $5.71\%$ $-32.91\%$
Oman $11.11\%$ $-45.68\%$
Philippines $13.77\%$ $-24.53\%$
Poland $5.06\%$ $-25.04\%$
Portugal $6.50\%$ $-20.45\%$
Romania $6.98\%$ $-13.46\%$
Russian Federation $9.41\%$ $-18.15\%$
Saudi Arabia $14.52\%$ $-36.71\%$
Serbia $4.96\%$ $-8.09\%$
South Africa $6.04\%$ $-24.29\%$
Spain $5.70\%$ $-27.76\%$
Sri Lanka $10.22\%$ $-5.08\%$
Sweden $4.24\%$ $-34.31\%$
Switzerland $5.93\%$ $-40.78\%$
Syrian Arab Republic $13.24\%$ $-24.92\%$
Tunisia $7.87\%$ $-44.33\%$
Turkey $8.89\%$ $-28.01\%$
United Arab Emirates $12.10\%$ $-44.14\%$
United Kingdom $4.97\%$ $-31.97\%$
United States $5.11\%$ $-29.73\%$
Zimbabwe $7.78\%$ $-21.09\%$
@@
  • [7] Whitney K. Newey and Kenneth D. West. A simple, positive semi-definite, heteroskedasticity and autocorrelation consistent covariance matrix. Econometrica, 55:703–708, 1987.

### v2

@@ Acknowledgments
Albania $12.26\%$ $31.89\%$
Algeria $10.30\%$ $46.19\%$
Australia $7.68\%$ $35.75\%$
Austria $9.53\%$ $47.00\%$
Bahrain $13.86\%$ $46.77\%$
Belarus $6.63\%$ $35.28\%$
Belgium $11.46\%$ $40.18\%$
Bosnia and Herzegovina $8.16\%$ $32.45\%$
Bulgaria $7.11\%$ $6.30\%$
Canada $7.45\%$ $40.34\%$
Cyprus $11.02\%$ $50.00\%$
Czechia $8.44\%$ $43.01\%$
Denmark $8.79\%$ $47.70\%$
Egypt $10.70\%$ $23.70\%$
Estonia $8.27\%$ $44.60\%$
France $9.11\%$ $34.78\%$
Germany $8.49\%$ $51.63\%$
Ghana $14.30\%$ $42.75\%$
Iceland $8.11\%$ $36.59\%$
Iran, Islamic Republic of $9.16\%$ $54.02\%$
Iraq $15.33\%$ $26.50\%$
Ireland $7.35\%$ $37.79\%$
Israel $8.85\%$ $40.69\%$
Italy $9.98\%$ $29.86\%$
Jamaica $4.95\%$ $14.81\%$
Japan $6.64\%$ $56.26\%$
Jordan $15.56\%$ $47.23\%$
Kenya $11.89\%$ $30.67\%$
Kuwait $10.43\%$ $39.39\%$
Lebanon $10.99\%$ $42.69\%$
Lithuania $9.99\%$ $32.43\%$
Moldova, Republic of $11.59\%$ $20.74\%$
Montenegro $7.54\%$ $35.03\%$
Morocco $14.08\%$ $38.63\%$
Netherlands $9.78\%$ $40.39\%$
New Zealand $6.90\%$ $42.72\%$
Nigeria $12.72\%$ $22.80\%$
Norway $8.49\%$ $38.37\%$
Oman $13.51\%$ $59.85\%$
Philippines $17.70\%$ $29.79\%$
Poland $8.14\%$ $30.25\%$
Portugal $8.44\%$ $23.57\%$
Romania $9.83\%$ $17.79\%$
Russian Federation $7.80\%$ $15.78\%$
Saudi Arabia $17.63\%$ $48.85\%$
Serbia $6.98\%$ $7.94\%$
South Africa $8.70\%$ $31.44\%$
Spain $8.66\%$ $34.10\%$
Sri Lanka $10.46\%$ $1.84\%$
Sweden $7.09\%$ $39.90\%$
Switzerland $9.09\%$ $46.99\%$
Syrian Arab Republic $15.17\%$ $33.67\%$
Tunisia $11.48\%$ $46.08\%$
Türkiye $12.07\%$ $35.74\%$
United Arab Emirates $15.68\%$ $59.48\%$
United Kingdom $7.18\%$ $40.19\%$
United States $7.44\%$ $37.03\%$
Zimbabwe $10.73\%$ $29.31\%$
Figure S1: Researcher population pyramids in 2010, 2023, and 2050 for Egypt, Tunisia, Japan, and the United States in Health Sciences. The number of active authors for each gender is displayed on a logarithmic horizontal axis. Total active author counts (NN) are provided for each panel. The 2050 pyramids and their corresponding counts ($\approx$) are projections based on 2023 trends.
Figure S2: Researcher population pyramids in 2010, 2023, and 2050 for Egypt, Tunisia, Japan, and the United States in Life Sciences. The number of active authors for each gender is displayed on a logarithmic horizontal axis. Total active author counts (NN) are provided for each panel. The 2050 pyramids and their corresponding counts ($\approx$) are projections based on 2023 trends.
Figure S3: Researcher population pyramids in 2010, 2023, and 2050 for Egypt, Tunisia, Japan, and the United States in Physical Sciences. The number of active authors for each gender is displayed on a logarithmic horizontal axis. Total active author counts (NN) are provided for each panel. The 2050 pyramids and their corresponding counts ($\approx$) are projections based on 2023 trends.
Figure S4: Researcher population pyramids in 2010, 2023, and 2050 for Egypt, Tunisia, Japan, and the United States in Social Sciences. The number of active authors for each gender is displayed on a logarithmic horizontal axis. Total active author counts (NN) are provided for each panel. The 2050 pyramids and their corresponding counts ($\approx$) are projections based on 2023 trends.
@@
  • [7] Louise Ackers. Internationalisation, mobility and metrics: A new form of indirect discrimination? Minerva, 46:411–435, 2008.